OpenStudy (anonymous):

find the volume of the solid generated by the following equations when rotated around the given axis: y=sqrtx y=0 and x=3 a)the y axis b)the line x=3 c)the line x=6

OpenStudy (anonymous):

yo amistre i need some help on some integrals as well... idk why i cant get them right

OpenStudy (amistre64):

we can shell it from 0 to 3 to rotate it about the x=0 axis

OpenStudy (anonymous):

cant shell, only washer and disk

OpenStudy (amistre64):

why cant you shell?

OpenStudy (anonymous):

not allowed to haha im not supposed to know how to shell yet

OpenStudy (amistre64):

do you gotta show your work? or just get the answer?

OpenStudy (anonymous):

well the thing is is that i can get the answer from the back, its the work that i cant get right for some reason

OpenStudy (amistre64):

to disk it around the x=0 or x=3 axis; we solve the eqs for y and intf(y)dy it

OpenStudy (anonymous):

are the limits also in terms of y? so 0 to sqrt3 ?

OpenStudy (amistre64):

yes; since you are traveling along the y axis to add up all the parts

OpenStudy (anonymous):

but i stillg et the wrong answer O.o

OpenStudy (amistre64):

y = 0 to y = sqrt(3)

OpenStudy (amistre64):

first figure the volume for the "cylinder" that is created by a radius of 3 and a height of sqrt(3)

OpenStudy (anonymous):

i use: \[\Pi \int\limits_{0}^{\sqrt{3}} (y ^{2})^{2}dy\]

OpenStudy (amistre64):

that is the area you wanna subtract yes ?

OpenStudy (anonymous):

subtract?

OpenStudy (anonymous):

i thought thats just the area i wanted XD

OpenStudy (anonymous):

oh i see. so i want to subtract it from the cylinder which is r=3 h=sqrt3?

OpenStudy (amistre64):

subtract the area created by the lower function from the area created by the upper function

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

your making a funny looking donut

OpenStudy (anonymous):

wait am i finding the area of the shaded, or the white?

OpenStudy (anonymous):

the shaded... right?

OpenStudy (amistre64):

the shaded is the area of the boundary; now we got to spin it about the x=0 axis; the x=6 axis and the x=3 axis for each part of the question

OpenStudy (amistre64):

the shaded area tells you what is spinning; rotate it to find the volume of revolution :)

OpenStudy (anonymous):

iight so for the x=3 and x=6... i cant get my euation straight ...

OpenStudy (amistre64):

why not: just subtract 3 and 6 respectively from the equations

OpenStudy (anonymous):

why subtract those?

OpenStudy (anonymous):

why not add or such?

OpenStudy (amistre64):

you want to get your axis of rotation to the origin its easier to calculate from that point

OpenStudy (amistre64):

sqrt(x-3) spins about the origin; sqrt(x-6) move the x = 6 axis to the origin etc..

OpenStudy (amistre64):

the heights never change :) do they

OpenStudy (anonymous):

ok i get it, idk why im getting so lost with concepts i used to know cold -.-

OpenStudy (amistre64):

keep in practice :)

OpenStudy (anonymous):

can you show me how to evaluate the integral of sech(x/2)

OpenStudy (anonymous):

need more practice.. If you get rusty, just find someone to tutor. That's what I do. ;)

OpenStudy (amistre64):

sech? cant say that Ive had to much practice with hyperbolics ...

OpenStudy (anonymous):

i come up with 4arctan(x/2) but then apparently because i have the area bounded by -4 and 4, it makes it 8arctan(x/2)-2pi... WHY!?!?!??!

OpenStudy (anonymous):

well im sure if you treat it as 2/(e^x + e^-x) you can do it XD

OpenStudy (amistre64):

wolframalpha it and see what they can shed some light on

OpenStudy (anonymous):

in this case, x is (x/2)

OpenStudy (anonymous):

wolframalpha didnt help understand why it did that tho

OpenStudy (anonymous):

polpak.. can you do it? XD

OpenStudy (anonymous):

sech is an even function. So \[\int_{-a}^a f(x)dx = 2\int_0^af(x)dx\] For the even function you come up with for the rotation.

OpenStudy (anonymous):

.... oh wow, what about the -2pi ?

OpenStudy (anonymous):

and for the solids of rotation, when done around the x=3 and x=6, do my bounds change? how?

OpenStudy (anonymous):

Sorry, someone called. just a sec.

OpenStudy (anonymous):

yeahh buddy

OpenStudy (anonymous):

Changing the bounds on x will only change the limits of your integration if you are still rotating about the x axis. If you are rotating about some other line, then things get tricky.

OpenStudy (anonymous):

explainnn please :O

OpenStudy (anonymous):

explain what?

OpenStudy (anonymous):

how the bounds change

OpenStudy (anonymous):

The best thing to use here is the cylindrical coordinates, since you need the x, y axes to find the area on 2-d and then sweep it across the axis of your interest to find the volume. Follow my next post to find the procedure of computing the area when rotated across y-axis (question a), the others can be similarly found. The answer is \[4\sqrt{3} \pi\]

OpenStudy (anonymous):

\[\int\limits_0^{2\pi} \int\limits_0^3 \int\limits_0^{\sqrt{x}} dy dx d \theta\] This is the volume when you rotate around y-axis. Carry out the calculations, I guess I did the calculations correct, the answer is as above I mentioned. Here is a drawing of the procedure.